OpenStudy (anonymous):

cansomeone plz help me... Simplify completely 12x+36/x^2-4x-21 and find the restrictions on the variable.

OpenStudy (anonymous):

@dumbcow @abb0t

OpenStudy (loser66):

factor numerator , denominator. simplify them the restriction is the value of x which makes denominator = 0. that means you have to solve for x ( from the denominator)

OpenStudy (anonymous):

i got x+3/x-7 \[x \neq7\],\[x \neq-3\]

OpenStudy (anonymous):

for my answer is that right

OpenStudy (loser66):

nope

OpenStudy (anonymous):

ok i will try again

OpenStudy (loser66):

\[\frac{12x+36}{x^2-4x-21}=\frac{12(x+3)}{(x+3)(x-7)}=\frac{12}{x-7}\] so restriction is x-7=0 or x = 7 the solution is just x \(\neq7\)